# Can LLMs Maintain Fundamental Abilities under KV Cache Compression?

### v1

@@ B.1 Hyper-parameters
MMLU (Hendrycks et al., 2020) 5 5 - - 0,5 - -
CommonsenseQA (Talmor et al., 2019) 4 4 - - - - -
GSM8K (Cobbe et al., 2021) 8 8 1-8 50 0,8 - -
HumanEval (Chen et al., 2021) 8 8 - - - - -
JailBreakV (Luo et al., 2024) 8 8 - - - - -
LongGenBench-GSM8K (Liu et al., 2024d) - - - - - 35 20
@@ B.2 Detail Results
KV cache compression methods show task-dependent performance degradation, with varying sensitivity thresholds across different benchmark categories.
@@
R1-GSM8K Baseline R1-Distill-Llama-8B FullKV: 0.6938
@@
GSM8K Baseline LLaMA-3.1-8B-Instruct FullKV: 0.7945
@@
Table 6: Performance Comparison of Different KV Cache Compression Methods on Multiple Benchmarks
@@
Table 7: KV Cache Compression Performance Comparison on GSM8K with Different Instruction TuningSettings
@@
Chunk-level compression is more effective for long-context arithmetic reasoning tasks. As shown in Table 9, ChunkKV demonstrates superior robustness across different compression ratios, particularly under aggressive compression settings. While other methods show significant performance degradation at 10% compression ratio (StreamingLLM: -9.8%, H2O: -37.8%, SnapKV: -17.1%, PyramidKV: -14.6%), ChunkKV maintains relatively stable performance with only a -3.7% drop. This stark contrast in performance suggests that chunk-level compression better preserves the essential contextual information needed for complex reasoning tasks. The method’s effectiveness likely stems from its ability to maintain the structural integrity of related context segments, which is particularly crucial for tasks requiring extended logical reasoning and arithmetic operations.
@@ C.1 Pseudocode
0:  Prompt with $n$ shots $\{s_{1},...,s_{n}\}$, prefill ratio $r_{p}$, decoding ratio $r_{d}$
0:  Compressed KV cache $KV_{\text{total}}$
1:  // Phase 1: Prefill Compression (performed once)
2:  for each shot $s_{i}$ in $\{s_{1},...,s_{n}\}$ do
3:     Compute $\text{Score}_{\text{prefill}}(s_{i})=\frac{1}{k_{i}}\sum_{t\in s_{i}}\sum_{h=1}^{H}\sum_{l=1}^{L}\alpha_{t,h}^{l}$
4:  end for
5:  Sort shots by $\text{Score}_{\text{prefill}}(s_{i})$ in descending order
6:  $S_{\text{preserved}}\leftarrow$ Select shots until $\sum_{s_{i}}k_{i}\leq r_{p}\times|KV_{\text{prefill}}|$
7:  $KV^{C}_{\text{prefill}}\leftarrow\text{Compress}(\{s_{i}|s_{i}\in S_{\text{preserved}}\})$
8:  // Phase 2: Decoding Compression (performed dynamically)
9:  for each decoding step do
10:     for each token $t$ in $KV_{\text{decoding}}$ do
11:        Compute $\text{Score}_{\text{decoding}}(t)=\sum_{h=1}^{H}\sum_{l=1}^{L}\alpha_{t,h}^{l}$
12:     end for
13:     $k\leftarrow r_{d}\times|KV_{\text{decoding}}|$
14:     $KV^{C}_{\text{decoding}}\leftarrow\text{TopK}(KV_{\text{decoding}},\text{Score}_{\text{decoding}},k)$
15:  end for
15:  $KV^{C}_{\text{prefill}}\cup KV^{C}_{\text{decoding}}$
@@ D.1 Dataset Details
MMLU (Hendrycks et al., 2020) World Knowledge 14,079 Accuracy Generation-Based
GSM8K (Cobbe et al., 2021) Arithmetic 1,319 Exact match Generation-Based
CSQA* (Talmor et al., 2019) Commonsense 1,221 Accuracy Generation-Based
HumanEval (Chen et al., 2021) Code Generation 164 Pass@1 rate Generation-Based
JailBreakV (Luo et al., 2024) Safety 28,000 Attack success rate Generation-Based

### v2

@@ B.1 Hyper-parameters
MMLU [33] 5 5 - - 0,5 - -
CommonsenseQA [35] 4 4 - - - - -
GSM8K [34] 8 8 1-8 50 0,8 - -
HumanEval [36] 8 8 - - - - -
JailBreakV [37] 8 8 - - - - -
LongGenBench-GSM8K [38] - - - - - 35 20
@@ B.2 Detail Results
KV cache compression methods show task-dependent performance degradation, WK and CSR are more robust to KV cache compression.
@@
R1-AR Baseline R1-Distill-Llama-8B FullKV: 0.6938
@@
AR Baseline LLaMA-3.1-8B-Instruct FullKV: 0.7945
@@
Table 7: Performance Comparison of Different KV Cache Compression Methods on KVFundaBench
@@
Table 8: KV Cache Compression Performance Comparison on Arithmetic Reasoning with Different Instruction TuningSettings
@@
Chunk-level compression is more effective for long-context structured reasoning tasks. As shown in Table 10, ChunkKV demonstrates superior robustness across different compression ratios, particularly under aggressive compression settings. While other methods show significant performance degradation at 10% compression ratio (StreamingLLM: -9.8%, H2O: -37.8%, SnapKV: -17.1%, PyramidKV: -14.6%), ChunkKV maintains relatively stable performance with only a -3.7% drop. This stark contrast in performance suggests that chunk-level compression better preserves the essential contextual information needed for complex reasoning tasks. The method’s effectiveness likely stems from its ability to maintain the structural integrity of related context segments, which is particularly crucial for tasks requiring extended logical reasoning and arithmetic operations.
@@
B.3 More experiments on other models
To further validate the generality of our findings, we also evaluate the impact of KV cache compression on a different model, Mistral-7B-Instruct. As shown in Figure 9, we observe that various KV cache compression methods lead to significant performance degradation across multiple fundamental tasks, especially under aggressive compression ratios. This result demonstrates that the reduction in foundation abilities due to KV cache compression is not limited to a specific model family, but is a general phenomenon affecting different LLM architectures.
Figure 9: Performance Comparison of KV Cache Compression Methods Across Tasks with Mistral-7B-Instruct.
@@ C.1 Pseudocode
1:Prompt with $n$ shots $\{s_{1},...,s_{n}\}$, prefill ratio $r_{p}$, decoding ratio $r_{d}$
2:Compressed KV cache $KV_{\text{total}}$
3:// Phase 1: Prefill Compression (performed once)
4:for each shot $s_{i}$ in $\{s_{1},...,s_{n}\}$ do
5:     Compute $\text{Score}_{\text{prefill}}(s_{i})=\frac{1}{k_{i}}\sum_{t\in s_{i}}\sum_{h=1}^{H}\sum_{l=1}^{L}\alpha_{t,h}^{l}$
6:end for
7:Sort shots by $\text{Score}_{\text{prefill}}(s_{i})$ in descending order
8:$S_{\text{preserved}}\leftarrow$ Select shots until $\sum_{s_{i}}k_{i}\leq r_{p}\times|KV_{\text{prefill}}|$
9:$KV^{C}_{\text{prefill}}\leftarrow\text{Compress}(\{s_{i}|s_{i}\in S_{\text{preserved}}\})$
10:// Phase 2: Decoding Compression (performed dynamically)
11:for each decoding step do
12:     for each token $t$ in $KV_{\text{decoding}}$ do
13:         Compute $\text{Score}_{\text{decoding}}(t)=\sum_{h=1}^{H}\sum_{l=1}^{L}\alpha_{t,h}^{l}$
14:     end for
15:     $k\leftarrow r_{d}\times|KV_{\text{decoding}}|$
16:     $KV^{C}_{\text{decoding}}\leftarrow\text{TopK}(KV_{\text{decoding}},\text{Score}_{\text{decoding}},k)$
17:end forreturn $KV^{C}_{\text{prefill}}\cup KV^{C}_{\text{decoding}}$
C.2 Latency and Throughput
We further compare the inference efficiency of ShotKV and the FullKV baseline in terms of latency and throughput under different input and output sequence lengths. As shown in Table 11, ShotKV consistently reduces latency and improves throughput compared to FullKV. For example, with an input length of 8192 and output length of 4096, ShotKV achieves an 11.3% reduction in latency and a 13.1% increase in throughput. These results demonstrate that ShotKV not only maintains model performance under aggressive KV cache compression, but also brings tangible efficiency benefits for long-context inference.
Table 11: Latency and throughput comparison between ShotKV and FullKV under different input-output configurations. Percentages in parentheses indicate improvements over FullKV baseline. The experiments test on the A40 server with batch size 1.
Method Sequence Length Performance Metrics
Input Output Latency(s) $\downarrow$ Throughput(T/S) $\uparrow$
FullKV 4096 4096 175.50 37.73
ShotKV 4096 4096 162.85 (7.2%) 41.12 (9.0%)
FullKV 8192 4096 183.42 55.93
ShotKV 8192 4096 162.78 (11.3%) 63.24 (13.1%)
@@ D.1 Dataset Details
MMLU [33] World Knowledge 14,079 Accuracy Generation-Based
GSM8K [34] Arithmetic 1,319 Exact match Generation-Based
CSQA [35] Commonsense 1,221 Accuracy Generation-Based
HumanEval [36] Code Generation 164 Pass@1 rate Generation-Based
JailBreakV [37] Safety 28,000 Attack success rate Generation-Based
LongGenBench [38] Long-Context Generation 23,000 Accuracy Generation-Based